OpenStudy (anonymous):

Given the parent functions f(x) = 5x - 1 and g(x) = 3x - 9, what is g(x) - f(x)

OpenStudy (snowfire):

Don't you just do 5x-1 - 3x-9? Or am I missing something?

OpenStudy (snowfire):

I mean, the other way around, but yeah

OpenStudy (snowfire):

3x-9-5x+1

OpenStudy (anonymous):

I really dont know caause I dont knoe how to do them but these are the answers a. g(x) - f(x) = 3x - 8 - 5x b. g(x) - f(x) = 3x - 10 - 5x c. g(x) - f(x) = -2x - 8 - 3x d. g(x) - f(x) = -2x - 10 - 3x

OpenStudy (anonymous):

its a i guess cuz i had a similiar question my teacher had told me some method i think its a

OpenStudy (anonymous):

no a is wrong

OpenStudy (anonymous):

3x-9-5x+1 will be the anser

OpenStudy (anonymous):

pm me

OpenStudy (anonymous):

is in it c ?

OpenStudy (snowfire):

You just combine like terms, but I guess the answer choices left out combining the x's for some reason.

OpenStudy (anonymous):

You wrote the second equation wrong, it is supposed to be 3^x - 9.